OpenStudy (anonymous):

What is this midpoint of the segment whose endpoints are (-6,5) and (4,-7)?

OpenStudy (ivettef365):

use formula for midpoint

sam (.sam.):

Given \((x_1,y_1)\), \((x_2,y_2)\) \[\text{Midpoint}~~M=(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})\] Substitute and solve

OpenStudy (anonymous):

\[\frac{ -2 }{ 2 }, \frac{ -2 }{2 } = ?\]